OpenStudy (anonymous):

Perform this operation and report the answer to the correct number of significant figures: What is the mass of a cube of aluminum that is 4.0 cm on each edge? The density of aluminum is 2.7 g/cm3. Show all calculations leading to an answer.help me please

OpenStudy (taramgrant0543664):

First step is to calculate the volume of the cube V= lengh x width x height We know each of these measurements are 4 cm so the volume is?

OpenStudy (taramgrant0543664):

Once you find the volume you can then use the formula M=VD where M is the mass that you're trying to find, V is the volume you found in the first step and D is the density that is given

OpenStudy (taramgrant0543664):

Rules when multiplying for significant figures are: a) First perform all the operations, even if changing from one formula to another. b) Round off the result so that it has the same number of sig figs as the least of all those used in your calculation.

OpenStudy (anonymous):

2.Vcube = length*width*height Vcube = 4.0cm*4.0cm*4.0cm Vcube = 64cm^3 2.7g/cm^3 = m/64cm^3 m = 2.7g/cm^3*64cm^3 m = 172.8g is hat right

OpenStudy (taramgrant0543664):

172.8 is the answer but that doesn't take sig figs into account

OpenStudy (anonymous):

okay im totally lost

OpenStudy (taramgrant0543664):

the two numbers have two significant digits therefore your answer can only have two significant figures

OpenStudy (anonymous):

sorry can you explain more

OpenStudy (joannablackwelder):

The numbers given in the problem (4.0 cm and 2.7 g/cm^3) each have 2 significant figures.

OpenStudy (joannablackwelder):

And since the problem used multiplication, and not addition, the answer must also have 2 significant figures. Does that make sense?

OpenStudy (anonymous):

littl

OpenStudy (taramgrant0543664):

If I had 6.00386 times 2.1 my answer would only have two significant digits because of the 2.1

OpenStudy (joannablackwelder):

Do you understand how to find the number of sig figs in a number?

OpenStudy (anonymous):

no

OpenStudy (taramgrant0543664):

Non-zero digits are always significant. Any zeros between two significant digits are significant. A final zero or trailing zeros in the decimal portion ONLY are significant.

OpenStudy (taramgrant0543664):

6 is one sig fig 606 is 3 sig figs 0.006 is 1 sig fig 0.0060 is 2 sig figs 60 is 1 sig fig

OpenStudy (joannablackwelder):

I use these two rules for finding the number of sig figs: When the decimal is present, start counting from the left side of the number with the first non-zero digit and count the rest of the digits. For example, 0.003040 has 4 sig figs.

OpenStudy (joannablackwelder):

When the decimal is absent, start counting from the right with the first non-zero digit. For example, 604,000 has 3 sig figs.